OpenStudy (anonymous):

Simplify square root of 2 multiplied by the cube root of 2 2 to the power of 1 over 6 2 to the power of 2 over 3 2 to the power of 5 over 6 2 to the power of 7 over 6

jagr2713 (jagr2713):

can you take a picture of it? or make it as an equation

OpenStudy (anonymous):

Sure

OpenStudy (anonymous):

that would really help

OpenStudy (anonymous):

\[(\sqrt{2}) (3\sqrt{2})\]

jagr2713 (jagr2713):

Ok first we simplify \[3\sqrt{2}\sqrt{2}\]

jagr2713 (jagr2713):

then we put them together to get \[3\sqrt{2\times 2}\] Do you know PEMDAS?

OpenStudy (anonymous):

Yes of course.

jagr2713 (jagr2713):

wait forget everything i said, its in wrong

OpenStudy (anonymous):

When I did it I got C for my answer.

jagr2713 (jagr2713):

How?

OpenStudy (anonymous):

Hold on BRB.